OpenStudy (anonymous):

Multiply , and simplify your answer. 8a/24* 8/20b

OpenStudy (anonymous):

8a/60b i think

OpenStudy (anonymous):

\[(8a/24) * (8/20b) = (8/24)(8/20)(a/b) = (1/3)(2/5)(a/b) = (2/15)(a/b) = 2a/15b\]